OpenStudy (anonymous):

4y-8-2y+5=0. When I did this I got 1.5.. is that right? If not, whatd I do wrong?

OpenStudy (jim766):

4y - 8 - 2y + 5 = 0 2y -3 = 0 2y = 3 y = 3/2 or y = 1.5 looks good!

OpenStudy (anonymous):

Alright thanks!(: